# A Multimodal Benchmark for Video Fake News Detection

### v1

@@ Abstract
The advent of multi-modal large language models (MLLMs) has greatly advanced research into applications for Video fake news detection (VFND) tasks. Traditional video-based FND benchmarks typically focus on the accuracy of the final decision, often failing to provide fine-grained assessments for the entire detection process, making the detection process a black box. Therefore, we introduce the MVFNDB (Multi-modal Video Fake News Detection Benchmark) based on the empirical analysis, which provides foundation for tasks definition. The benchmark comprises 10 tasks and is meticulously crafted to probe MLLMs’ perception, understanding, and reasoning capacities during detection, featuring 9,730 human-annotated video-related questions based on a carefully constructed taxonomy ability of VFND. To validate the impact of combining multiple features on the final results, we design a novel framework named MVFND-CoT, which incorporates both creator-added content and original shooting footage reasoning. Building upon the benchmark, we conduct an in-depth analysis of the deeper factors influencing accuracy, including video processing strategies and the alignment between video features and model capabilities. We believe this benchmark will lay a solid foundation for future evaluations and advancements of MLLMs in the domain of video fake news detection.
Perception, Understanding and Reasoning,
@@ A Multimodal Benchmark for Video Fake News Detection
Yakun Cui1, Fushuo Huo2, Weijie Shi1, Juntao Dai4, Hang Du3, Zhenghao Zhu1, Sirui Han1 Yike Guo1 1The Hong Kong University of Science and Technology 2The Hong Kong Polytechnic University 3Beijing University of Posts and Telecommunications 4Peking University ycuibd@connect.ust.hk {siruihan, yikeguo}@ust.hk
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success in tackling a wide range of general AI tasks in recent years. However, certain tasks may require multi-modal capabilities to process visual and textual information and obtain better performance. To address these needs, MLLMs have been developed and widely adopted for more complex, multi-modal tasks. With the advancement of MLLMs, numerous benchmarks have been proposed to evaluate their capacities when processing multi-modal information. These assessments typically focus on general problem-solving abilities, such as video captioning, grounding, and object counting; However, due to expert knowledge requirements, task-specific scenarios, such as video-based FND, remain under explored. Firstly, the majority of previous video-based FND benchmarks were primarily designed for classification models. However, the mechanism of these benchmarks for VFND tasks differs fundamentally from that of MLLMs in processing VFND. Therefore, proposing a novel benchmark that aligns with the paradigm of MLLMs can evaluate the cross-modal understanding ability and knowledge generalization capability of MLLMs, and pave the way for further exploration of MLLMs in this field. Secondly, lack of an assessment of interpretable results. Existing benchmarks for video-based VFND focus exclusively on the distribution of classification results. However, the detection process is practically not a black-box, it requires evidence-based inference generation. The lack of evaluation of interpretable outcomes prevents us from tracing back process of detection and deprives us of direct access to more targeted optimization guidance. Thirdly, clues for VFND lie in multiple features of the video. Consequently, VFND is a process-and-result-oriented task. However,the majority of existing benchmarks predominantly focus on result-oriented evaluation; they lack the design of corresponding tasks to assess these specific capabilities of MLLMs and the extent to which these capabilities influence the final detection results. This limitation prevents us from optimizing FND performance based on in-depth analyses of the multimodal capabilities inherent in MLLMs. Consequently, there is a pressing need to establish a more comprehensive and targeted benchmark tailored to these specialized requirements. Given the mentioned gap, the benchmark presents the following critical advantages:
1) Comprehensive evaluation of both open-source and proprietary MLLMs on MVFND. Specifically, we evaluated the performance of both open-source and proprietary MLLMs across 10 tasks on the new test dataset, which features a more balanced news category and label distribution. Meanwhile, the selected models include both the sota models in the current video processing field and the most widely used open-source models. To explore the impact of model scale on performance, we also incorporated 3 versions of the same-type model with different scales. Our work serves as a dedicated benchmark for MLLMs in the field of MVFND.
2) Process-oriented evaluation for MVFND. We propose 10 distinct tasks that collectively evaluate the perception, understanding, and reasoning capabilities of MLLMs for MVFND process and results. On one hand, these tasks not only include general video-related tasks but also incorporate specialized tasks that are , and they ensure that the evaluation covers the entire detection process. On the other hand, for the metrics, the alignment between task properties and metric formats enables both tasks with high accuracy requirements (e.g., counting, recognition) and tasks with high semantic matching requirements (e.g., reasoning) to be effectively evaluated.
3) Applicable evaluation tasks for MVFND. Each task in the benchmark is defined based on the differentiated features of real and fake video news—these features are identified through empirical analysis. For one thing, the source video data come from original content on social platforms. This helps reduce bias in these features. For another, these features match real-world situations, and some have been validated by authoritative studies in the news field. Thus, these tasks should theoretically have certain general applicability in the MVFND.
4) Further Exploration based on MVFNDB. For one thing, we proposed a framework——MVFND-COT, which derived from evaluation result and verify the effectiveness of combination of creator-added text and original shooting footage reasoning. For another, the process-based evaluation enables us to identify the impacts of video processing strategies and non-visual features (e.g., duration, category) on VFND. Furthermore, this evaluation leads to the first discovery of a performance bottleneck that arises when there is a mismatch between model capabilities and video features.
@@ 2 Related Work
2.1 Micro-Video Fake News Detection
Fake News Detection (FND) seeks to evaluate the veracity of news by analyzing a range of characteristics across multiple modalities Wang et al. (2024b); D’ulizia et al. (2021); Qi et al. (2023); Papadopoulou et al. (2019), including text, images, audio, and video. Early research in this field Cao et al. (2020); Bu et al. (2024) primarily concentrated on extracting discriminative features from textual data. For instance, Wei et al. (2025) proposed a dual-granularity adversarial training approach that integrates document-level and entity-level features Wu et al. (2024); Shen et al. (2025), enabling models to better adapt to unseen target domains.
Compared to textual formats, videos inherently convey richer information because they combine multiple data types and complex cross-modal interactions Zong et al. (2024), Zheng et al. (2022), Wu et al. (2023). For example, Qi et al. (2023) demonstrated that integrating video content, user comments, and textual metadata (such as titles) collected from platforms like DouYin and KuaiShou can significantly enhance the effectiveness of fake news verification.
2.2 Multi-modal Perception, Understanding and Reasoning over Video
The emergence of extensive multi-modal datasets Yuan et al. (2025), alongside architectural innovations, has significantly propelled the development of multi-modal large language models (MLLMs). These advancements have enabled MLLMs to achieve competitive performance in core multi-modal tasks such as cross-modal perception Tang et al. (2025a), object grounding Wasim et al. (2024), and image captioning Bucciarelli et al. (2024). As a result, there has been growing research interest in exploring the multi-modal reasoning capabilities of MLLMs, exemplified by approaches such as multi-modal Chain-of-Thought (CoT) Wang et al. (2025b) and VisualPRM Wang et al. (2025a).
Videos inherently possess dynamic temporal features—including motion patterns Qian et al. (2024b); Zhao et al. (2023); Wang et al. (2024c); Maaz et al. (2023), audio rhythms, and color transitions—that go beyond what static text and images can offer, providing unique signatures crucial for analysis. For instance, Video-of-Thought Fei et al. (2024) introduced spatial-temporal grounding-aware tuning, which effectively bridges perceptual insights with cognitive reasoning. Video-R1 Feng et al. (2025) further advanced reasoning capabilities, achieving superior performance across diverse video benchmarks compared to previous methods. Additionally, Momentor Qian et al. (2024a) performed segment-level reasoning and localization, while CaRDiff Tang et al. (2025b) utilized a three-stage reasoning process to reach state-of-the-art results in video saliency prediction.
2.3 MLLM Benchmark over Video
As MLLMs continue to demonstrate increasingly robust multi-modal capabilities, a variety of benchmarking frameworks Ning et al. (2023); Zhang et al. (2025); Caba Heilbron et al. (2015); Liu et al. (2024) have been developed to evaluate their performance in perception, understanding, and reasoning. For example, MMMU Yue et al. (2024) provides a systematic framework for assessing visual perception and reasoning abilities across a broad range of disciplines, including the arts and sciences.
MMBench-Video Fang et al. (2024) incorporates long-form videos to assess MLLMs’ spatial-temporal understanding, while MLVU Zhou et al. (2025) evaluates models on tasks such as topic reasoning, anomaly recognition, and summarization. However, most existing benchmarks focus on generic multi-modal capabilities Han et al. (2025); Sun et al. (2025); Wang et al. (2024a), often overlooking specialized domain-specific tasks that demand nuanced and targeted reasoning. Fake news detection uniquely requires theme-dependent verification, involving analyses that range from low-level feature integrity (e.g., HSV anomalies) to high-level semantic coherence. Unlike general benchmarks, our proposed benchmark is designed to specifically evaluate MLLMs’ discriminative perception, thematic understanding, and evidential reasoning skills, using veracity-relevant feature differentiators as the basis for assessment.
3 Empirical Analysis
Our benchmark aims to comprehensively evaluate MLLMs’ capabilities in both detection result and process of MVFND. Higher accuracy performance of the model can only be achieved when it accurately perceives the key features in the video during the recognition process. Therefore, we conduct an empirical analysis on a real-world dataset——FakeSV, to analyze and demonstrate the differences between real and fake news in terms of several multi-modal characteristics. Additionally, we provide plausible explanations for these phenomena from the perspectives of the purpose of news productionFirmstone (2024) and the background of the producers. We analyses the creator-added content and original shooting footage respectively, as the former, with lower credibility, often embeds a higher degree of the author’s creative intent Guo et al. (2025). We only demonstrate some key findings from FakeSV here due to space limitations and attach remain results in the Appendix A.
3.1 Creator-added Content
Creator-added text is typically used in video news to highlight key information, such as sites, date. We employee the PaddleOCR tools to extract text regions from video and process them with the pipeline in Figure 10.
Figure 1: Hue distribution in creator-added content
of fake/real news video
3.1.1 Color Distribution
Finding. In terms of font color selection, fake news typically utilize colors with hue values within the range 0°-5°, whereas real news producers prefer to colors ranging 25°-30°.
Analysis. Given the prior work Zhou et al. (2022) has indicated font color can affect the understanding ability of news readers and Hue acts as the primary identifier of a color, we conduct a census comparing the pixel-level distribution of hues in real and fake video’s creator-added text. As demonstrated in Figure 1, the results indicate a notable divergence in hue distribution between real and fake news videos. We attribute this differentiated distribution to the fake news producers tend to utilize high emotional color to influence audiences’ cognition, however, real news creators employee formal color to guide users to focus on key information.
(a) Distribution of text region in fake news video
(b) Distribution of text region in real news video
Figure 2: Comparison of heatmaps
3.1.2 Spatial Distribution
Finding. In terms of text position arrangement, the text positions in fake news exhibit a higher degree of randomness, whereas those in real news show a more concentrated distribution.
Analysis. The placement selection of text shall not only ensure that the audience can effectively perceive the prompt information, but also minimize occlusion of the original shooting video frames—a balance that is critical to maintaining both the information delivery efficiency and the visual integrity of the news content. Our analysis focus on the text’s spatial distribution on screen. Figure 2 highlights the distinctions between fake and real news videos. We ascribe this discrepancy to the following factors, 1) fake news producers often overlay text on video footage to obscure factual details in the original visuals. 2) fake news producers lack professional creation training about screen-text arrangement.
3.2 Original Shooting Footage
Typically, original footage in news videos conveys dynamic information about reported events. To investigate differences in original footage between fake and real news videos, we focus on analyzing consecutive video frames. These findings provide evidence supporting the tasks in our benchmark, designed to evaluate the ability of MLLMs to perceive dynamic information.
3.2.1 Key footage Distribution
Figure 3: Distribution of key footage in fake/real news video
Finding In terms of on-site shooting footage, real news contains a relatively higher proportion of such content across all time segments. Regarding close-up shooting footage of characters, real news tends to arrange such shots in the tail section of the video, whereas fake news prefer to place them at the beginning. As for official declarations, real news incorporates more these footage in the tail part of the video.
Analysis Key frames are closely tied to the news theme and hold significant evidentiary value for veracity detection. To examine the temporal distribution of key footage, we analyz the proportion of three type of key footage——on-site shooting, close-up shooting and official declarations across eight temporally sequential video segments. Figure 3 illustrates how the proportion of key footage in real and fake news videos varys over time.
4 The MVFNDB Benchmark
4.1 Overview of MVFNDB
We propose MVFNDB, which builds upon the video clips of the open-source FakeSV dataset Qi et al. (2023). MVFNDB is a process and result oriented benchmark consisting 10 tasks evaluating MLLMs’ perception, understanding and reasoning capacities during FND. The test set comprises 10.5k questions, which cover 3 distinct formats: single-choice, multiple-choice, and open-ended generation. The metrics include both exact match and semantic match. The benchmark is distinguished by the following features:
Extensive Evaluation Datasets.Videos in FakeSV are collected from real-world platforms, so this dataset enables our evaluation to align as closely as possible with the prevalent distribution of fake news in terms of content styles, editing characteristics and misinformation creation patterns.
Diversified Evaluation Tasks. MVFNDB provides a diverse array of evaluation tasks, which are closely aligned with perception, understanding and reasoning capacities of MLLMs, such as temporal grounding, objects counting, video captioning. All tasks are tailored for MVFND process, that is to say, each task is associated with the final detection result. Specifically, the better a model performs on a single individual task, the higher the accuracy of its FND should be. The tasks encompass three categories: single-choice, multi-choice, and free-form generation. For each task category, its format is strictly tailored to align with the inherent attributes and requirements of the task type. This design enables the benchmark to assess MLLMs’ both video contents perception and reasoning semantic expression capacities.
Comprehensive Process and Result Evaluation. To evaluate the end-to-end capabilities of MLLMs in FND, the MVFNDB uses the accuracy of binary classification results as a key metric. At the same time, it assesses the process of generating these results. This assessment includes evaluation of the final interpretable answer and 14 intermediate capabilities related to the detection process. The latter focuses on 3 core aspects of MLLMs in individual task: 1) Whether models can accurately identify important multi-modal features in videos based on task goals; 2) Whether they can correctly understand the main content and theme of the news; 3) Whether they can use appropriate general knowledge and video relevant characteristic to support reasoning process.
@@ 4.1 Overview of MVFNDB
Table 1: Details for Tasks and Datasets. Avg EHR: the average elements hit rate. Avg AD: the average absolute distance. Avg ENHR: the average entities hit rate. Avg RHR: the average rationale hit rate. Avg KHR: average knowledge hit rate. FC: the factual consistency. TR: the theme relevance. CO: Completeness. CAC: the creator-added content. OSF: the original shooting footage.
Task Task Format (output) Dataset(Train/Test) Category Target Feature Evaluation metrics
Key Elements Perception (KEP) Open-Ended with Specific Format (text) 2.87k/0.75k Perception OSF Avg EHR
Distinguish CAC and OSF (DCS) Open-Ended with Specific Format (text) 2.87k/0.75k Perception CAC and OSF ROUGE-L
Creator-added Text Color Perception(CCP) Single Selection (text) 5.46k/1.15k Perception CAC Accuracy
Creator-added Text 2D-position Perception (CPP) Single Selection (text) 5.19k/1.08k Perception CAC Accuracy
Key Elements Temporal Grounding (KEG) Open-Ended with Specific Format (digit, text) 2.87k/0.75k Perception OSF IoU
Shooting Angles Counting (SAC) Open-Ended with Specific Format (digit) 2.87k/0.75k Perception OSF Avg AD
Human Identity Recognition (HIR) Multi Selection (options) 2.87k/0.75k Perception OSF Accuracy
News Video Type Understanding (NTU) Open-Ended (text) 2.87k/0.75k Understanding CAC and OSF Accuracy
News Video Theme Understanding (NEU) Single Selection (option) 2.87k/0.75k Understanding CAC and OSF FC,TR,CO
Final Detection Reasoning (FDR) Open-Ended (text) 2.87k/0.75k Reasoning CAC and OSF Avg ENHR
Final Detection Reasoning (FDR) Open-Ended (text) 2.87k/0.75k Reasoning External Knowledge Avg KHR
Final Detection Reasoning (FDR) Open-Ended (text) 2.87k/0.75k Reasoning Rationale Avg RHR
Final Detection Reasoning (FDR) Open-Ended (text) / Reasoning Real/Fake Accuracy
4.2 Data curation
The videos annotated in our benchmark are sourced from the FakeSV dataset, which is curated by extracting videos directly from two prevalent social media platforms——Douyin and Kuaishou. Using this dataset gives our benchmark several advantages: 1) The original labels have been cross-validated against official rumor-refutation platforms, which ensures the factual accuracy and reliability of our annotations. 2) The platform-native nature of the FakeSV dataset make the videos retain the authentic characteristics of real-world social media content style, which endows our benchmark captures generalized features reflective of actual fake news creation and achieves close alignment with the practical realities of FND. To comprehensively evaluate MLLMs’ performance in MVFND, we constructed a novel test split dataset with a balanced label and category distribution. We classify news videos into six distinct categories, whose detailed descriptions are provided in Appendix B.1 and 20. This dataset enables a thorough assessment of MLLMs’ ability to handle various content categories. The test split comprises 10 tasks, each containing 753 pairs.
For data annotation, as depicted in Figure 4, to reduce bias and hallucination, we utilize different types of MLLMs (Qwen2.5vl-72b and Gemini2.5-flash) with strong video process capabilities to generate and refine visual tasks. Then, we use a LLM(GPT-o1) with excellent reasoning capabilities to evaluate and filter the reasoning outputs of MLLMs. The GPT-o1 model evaluates these reasoning outputs for factual consistency and logical rationality, assigning scores from 0 to 5. Outputs scoring below 4 in either metric are reprocessed by the Gemini. Finally, perception generations are verified by 20 news creation experts after 10 hours annotation training; understanding and reasoning outputs are reviewed and revised by 10 detection experts (with 5+ years of video news censor experience in online public opinion regulatory authority) and 5 domain experts(with 5+ years of specialized field research experience). This iterative process continues until all task data is fully processed.

### v2

# Perception, Understanding and Reasoning: A Multimodal Benchmark for Video Fake News Detection

@@ Abstract
The advent of multi-modal large language models (MLLMs) has greatly advanced research on video fake news detection (VFND) tasks. Existing benchmarks typically focus on the detection accuracy, while failing to provide fine-grained assessments for the entire detection process. To address these limitations, we introduce POVFNDB (Process-oriented Video Fake News Detection Benchmark), a process-oriented benchmark comprising 10 tasks designed to systematically evaluate MLLMs’ perception, understanding, and reasoning capabilities in VFND. This benchmark contains 36,240 human-annotated question-answer (QA) in structured or open-ended formats, spanning 15 distinct evaluation dimensions that characterize different aspects of the video fake news detection process. Using POVFNDB, we conduct comprehensive evaluations on both proprietary and open-source MLLMs. Moreover, we establish a strong benchmark baseline by fine-tuning Qwen2.5VL-7B-Instruct on process-oriented chain-of-thought data constructed with our proposed POVFND-CoT framework, achieving state-of-the-art performance on VFND.
Perception, Understanding and Reasoning:
@@ A Multimodal Benchmark for Video Fake News Detection
Yakun Cui1, Peng Qi2, Fushuo Huo3, Weijie Shi1, Juntao Dai5, Hang Du4, Zhenghao Zhu1, Sirui Han1 Yike Guo1 1The Hong Kong University of Science and Technology 2National University of Singapore 3The Hong Kong Polytechnic University 4Beijing University of Posts and Telecommunications 5Peking University ycuibd@connect.ust.hk {siruihan, yikeguo}@ust.hk
@@ 1 Introduction
With the rise of social media, video has emerged as an important medium for news dissemination. However, this shift facilitates the spread of fake news due to the high pervasion of video fake news Nan et al. (2024); Qi et al. (2023b); D’ulizia et al. (2021), posing risks to social stability and necessitating effective video fake news detection (VFND).
Most existing detection paradigms Zong et al. (2024); Qi et al. (2023a); Bu et al. (2024); Shen et al. (2025) are result-centric, focusing primarily on final detection accuracy rather than the intermediate reasoning process, as illustrated in Figure 1. However, fake news detection is inherently an evidence-based reasoning task rather than a simple black-box classification problem. It requires models to perform step-by-step reasoning grounded in evidence drawn from both external knowledge sources and news content to produce interpretable predictions. As a result, explainable fake news detection has recently emerged as an important research direction.
Figure 1: Comparison between result-centric and process-oriented evaluation. The MLLM correctly identifies entities (red), retrieves knowledge (blue), applies rationales (green), and reaches conclusions (orange), with all corresponding metrics (matching font colors) improving.
Compared to text-based fake news, video fake news is more challenging to model and evaluate, as videos convey information through multiple modalities with higher information density and complex temporal structures. Misleading cues can arise not only from semantic content but also from editing patterns and cross-modal interactions. Indeed, video fake news detection is a multi-faceted task that requires diverse capabilities beyond final veracity prediction. At the perception level, models must accurately recognize and localize key elements in videos, and distinguish between intrinsic video content and creator-added elements, such as overlaid text and shooting transitions, which often play a critical role in misleading narratives. At a higher level of understanding and reasoning, VFND further requires models to comprehend the underlying news events and perform evidence-based factual reasoning by integrating external knowledge with video content. These requirements naturally align with the strengths of MLLMs, which leverage extensive world knowledge, multimodal perception and understanding, and semantic reasoning to generate coherent explanations. However, existing benchmarks provide only final veracity labels for classification Qi et al. (2023a); Wang et al. (2025c). The absence of supervision signals for intermediate steps prevents fine-grained assessment of the distinct capabilities involved in the VFND process, making it difficult to systematically analyze and optimize MLLMs for this task.
Figure 2: Framework of POVFNDB: Ten Evaluation Tasks Across Perception, Understanding and Reasoning.
Dataset Cls Per Und Rea #M Annt (Fake/Real)
CHECKED $\times$ $\times$ $\times$ 1 1.8k/0.34k
FakeSV $\times$ $\times$ $\times$ 1 1.8k/1.8k
TRUE $\times$ $\times$ 1 1.1k/1.8k
FakeVV $\times$ $\times$ 4 5.1k/5.1k
FMNV $\times$ $\times$ $\times$ 3 1.5k/0.9k
[…]
Table 1: Comparison of datasets for VFND. Cls: Classification, Per: Perception, Und: Understanding, Rea: Reasoning, #M: Number of metrics, Annt: Annotations.
To address these challenges, we develop POVFNDB, a process-oriented benchmark that supports fine-grained assessment of the diverse capabilities involved in video fake news detection. Built upon 3,624 videos from FakeSV Qi et al. (2023a), a widely used real-world video fake news dataset, POVFNDB constructs approximately 36,240 question–answer pairs organized into three core dimensions of perception, understanding, and reasoning, covering 10 sub-tasks and 15 evaluation metrics. Based on POVFNDB, we perform a thorough evaluation of mainstream MLLMs, including open-source and proprietary video or image MLLMs. Extensive experiments demonstrate that even the leading MLLM (i.e., Gemini2.5-Flash) exhibit subpar performance in most perception and reasoning sub-tasks, underscoring the challenges and necessity of multimodal research tailored for the VFND process. Motivated by insights from rationale validation, we fine-tune Qwen2.5VL-7B-Instruct on a reasoning dataset generated by our proposed POVFND-CoT, a chain-of-thought method. This approach achieves state-of-the-art performance, highlighting the importance of developing domain-specific reasoning strategies for VFND. Our contributions can be summarized as follows:
  • Novel Process-Oriented Benchmark: The proposed POVFNDB benchmark is organized around three core dimensions, perception, understanding, and reasoning, and comprises 10 sub-tasks with 15 evaluation metrics based on 36,240 question–answer annotations. This design enables fine-grained evaluation and diagnostic analysis of model capabilities throughout the video fake news detection pipeline.
  • Comprehensive Evaluation of MLLMs: Using POVFNDB, we extensively evaluate mainstream MLLMs, including open-source/proprietary and image-/video-based models, revealing their strengths and limitations across VFND sub-tasks and diverse capability dimensions.
  • Strong Explainable Detection Baseline: We further propose POVFND-CoT, an explainable chain-of-thought framework for VFND that leverages process-aware reasoning signals. By fine-tuning models on reasoning data generated under POVFND-CoT, we establish a strong baseline that achieves state-of-the-art performance on video fake news detection.
@@ 2 Related Work
2.1 Video Fake News Detection
Fake News Detection (FND) seeks to evaluate the veracity of news by analyzing characteristics across multiple modalities Xu and Yan (2025); Zhang et al. (2025a); Guo et al. (2025); Wang et al. (2024b), including text, images, audio, and video. Zhou and Zafarani (2019)leveraged network-based features and propagation patterns to enhance fake news detection for textual articles. SnifferQi et al. (2024) leverages MLLMs to provide explainable misinformation detection for image-based news. Video news convey richer informationWang et al. (2025a); Zong et al. (2025); Zeng et al. (2025); Wu et al. (2023).Qi et al. (2023a) demonstrated that integrating video content, comments, and metadata collected from social media platforms significantly enhance the effectiveness of VFND.
2.2 Video Benchmark for MLLM Evaluation
The emergence of extensive multi-modal datasets Yuan et al. (2025) and architectural innovations have significantly propelled the development of MLLMs. These advancements have enabled MLLMs to achieve competitive performance in core multi-modal tasks Tang et al. (2025a); Wang et al. (2025b). Videos exhibit inherent temporal dynamics, such as motion patterns Qian et al. (2024b); Wang et al. (2024c) and color transitions, offering unique analytical features that distinguish them from static text and image modalities. There has been growing interest in exploring MLLMs’ capacities for video-based tasks Qian et al. (2024a); Tang et al. (2025b); Feng et al. (2025). Video-of-Thought Fei et al. (2024) introduced spatial-temporal grounding-aware tuning, which effectively bridges perceptual insights with cognitive reasoning. As MLLMs continue to demonstrate robust multi-modal video processing performance, a variety of benchmarking Ning et al. (2023); Zhang et al. (2025b); Caba Heilbron et al. (2015); Liu et al. (2024) have been developed to evaluate their perception, understanding, and reasoning capacities. MMMU Yue et al. (2024) evaluates visual perception and reasoning abilities across a broad range of disciplines. MMBench-Video Fang et al. (2024) incorporates long-form videos to assess MLLMs’ spatial-temporal understanding.
However, these benchmarks focus on general-purpose multi-modal capabilities Han et al. (2025); Sun et al. (2025); Wang et al. (2024a), overlooking specialized domain tasks. As a comparison, our proposed benchmark is designed to specifically evaluate MLLMs’ discriminative visual perception, content understanding, and veracity reasoning abilities in the video fake news detection task.
3 Proposed Benchemark
Task Type Task Format Target Metrics
Perception Key Elements Perception (KEP) Open-Ended with format OSF Avg.EHR
Distinguish CAC and OSF (DCS) Open-Ended with format CAC&OSF ROUGE-L
Creator-added Text Color (CCP) Single Selection CAC Accuracy
Creator-added Text Position (CPP) Single Selection CAC Accuracy
Key Elements Grounding (KEG) Open-Ended with format OSF IoU
Shooting Angles Counting (SAC) Open-Ended with format OSF Avg.AD
Human Identity Recognition (HIR) Multi Selection OSF Accuracy
Understanding News Video Type (NTU) Open-Ended CAC&OSF Accuracy.
News Video Theme (NEU) Single Selection CAC&OSF Avg.FC/TR/CO
Reasoning Final Detection Reasoning (FDR) Open-Ended CAC&OSF Avg.ENHR
Ext.Know. Avg.KHR
Rationale Avg.RHR
Real/Fake Accuracy
@@ 3 Proposed Benchemark
Table 2: Task Description. CAC: Creator-added content. OSF: Original shooting footage. Ext.Know.: External Knowledge. Avg.EHR: Average elements hit rate. Avg.AD: Average absolute distance. Avg.ENHR: Average entities hit rate. Avg.RHR: Average rationale hit rate. Avg.KHR: Average knowledge hit rate. FC: Factual consistency. TR: Theme relevance. CO: Completeness.
In this section, we detail the construction of POVFNDB. We first introduce the definitions and motivations of the process tasks, then describe video selection and annotation procedures, present the resulting data distributions, and finally outline the task-specific evaluation methodology.